# FP-Inconsistent

### v1

@@ Abstract
As browser fingerprinting is increasingly being used for bot detection, bots have started altering their fingerprints for evasion. We conduct the first large-scale evaluation of evasive bots to investigate whether and how altering fingerprints helps bots evade detection. To systematically investigate evasive bots, we deploy a honey site incorporating two anti-bot services (DataDome and BotD) and solicit bot traffic from 20 different bot services that purport to sell “realistic and undetectable traffic.” Across half a million requests from 20 different bot services on our honey site, we find an average evasion rate of 52.93% against DataDome and 44.56% evasion rate against BotD. Our comparison of fingerprint attributes from bot services that evade each anti-bot service individually as well as bot services that evade both shows that bot services indeed alter different browser fingerprint attributes for evasion. Further, our analysis reveals the presence of inconsistent fingerprint attributes in evasive bots. Given evasive bots seem to have difficulty in ensuring consistency in their fingerprint attributes, we propose a data-driven approach to discover rules to detect such inconsistencies across space (two attributes in a given browser fingerprint) and time (a single attribute at two different points in time). These rules, which can be readily deployed by anti-bot services, reduce the evasion rate of evasive bots against DataDome and BotD by 48.11% and 44.95% respectively.
1 Introduction
@@
The proportion of bots on the web is on the rise [1]. As of 2023, bots constitute around 47.5% of online traffic [2], with 63.6% of those being bots that engage in malicious activity. Fraudsters employ such bots to launch a multitude of attacks [3, 4, 5, 6, 7] which results in of dollars of loss to the industry. To counter such attacks, anti-bot services aim to detect and block bot traffic. Prior research has shown that anti-bot services employ browser fingerprinting to detect bots without disrupting the user experience of legitimate users [8, 9]. Browser fingerprints capture attributes of the web browser sending web requests and anti-bot services attempt to use differences in these attributes to distinguish bots from real users [10].
@@
Blackhat marketplaces [11, 12, 13], however, advertise realistic and undetectable bot traffic as a service. The traffic from such services constitute impression fraud and serve to artificially boost website engagement for monetization [3, 14, 15]. We surmise that the bots from these services are likely manipulating their browser attributes, which are used by anti-bot services for detection, to alter their fingerprints [16, 17] for evasion. We refer to such bots as evasive bots and their fingerprints as evasive fingerprints. It is imperative to characterize evasive bots and their fingerprints to devise countermeasures against them to bolster bot detection.
@@
Prior research has studied bot fingerprints by employing their own bots [18, 8] or by focusing on naturally discovered bots on their honey sites [19]. Thus, this work does not capture the evasive fingerprints used by bots seeking to evade detection in the wild. Wu et al. performed a large scale characterization of the differences between human and bot fingerprints in the wild [10]. However, they did not characterize evasive bots since they treat their bot detection system decisions as ground-truth to distinguish between the fingerprints of bots and real users. Thus, they are unable to identify evasive bots that evade their bot detection system.
@@
To fill this gap, we perform the first large-scale measurement of evasive bots that are able to evade anti-bot services. To ensure reliable ground-truth, we use a honey site architecture to selectively drive traffic from different bot services from blackhat marketplaces to different instances of our honey site. Our honey site is designed such that the requests recorded at each instance is mapped to the corresponding bot service from whom we purchased traffic. These operators advertise their traffic as being realistic and natural, indicating that they likely employ evasive bots to ensure that they do not get detected. We integrate two commercial bot detection services (DataDome and BotD) on our honey site for bot detection.
@@
We receive 507,080 requests from 20 different bot services, with DataDome and BotD detected 55.44% and 47.07% of these requests respectively. We systematically analyze fingerprint attributes from different bot services to identify different sets of attributes that are effective at evading DataDome and BotD individually as well as attributes that are effective at evading both anti-bot services. Our analysis reveals spatial inconsistencies (among the different attributes of a given request) and temporal inconsistencies (across requests originating from the same device). These were egregious inconsistencies that cannot exist for real users, thereby making them outright signatures to detect bots.
@@
We use insights from our analysis to develop FP-Inconsistent, a data-driven, semi-automatic approach to discover inconsistencies in browser fingerprint attributes for more reliable bot detection. FP-Inconsistent relies on the fundamental understanding that real devices can only have limited number of hardware and software configurations that are reflected in browser fingerprint attributes. Evasive bots, in their attempt to evade detection, proliferate large number of invalid or extraneous configurations. FP-Inconsistent leverages this mismatch between expected and observed number of configurations to identify potential inconsistencies among evasive bot fingerprints. It does so by calculating the number of configurations for pairs of evasive bot fingerprint attributes, and identifying inconsistencies among attribute pairs that exhibit higher than expected number of configurations. Using this approach, we generate inconsistency rules that can be readily deployed by anti-bot services. While prior research has proposed the use of inconsistencies for bot detection [20, 8], it mainly relies on one-off anecdotes to define inconsistencies which does not scale because it is not data-driven. FP-Inconsistent is the first technique to systematize the generation of inconsistency based rules for bot detection. Our evaluation shows the rules generated by FP-Inconsistent are able to achieve 48.11% and 44.95% reduction in requests that evade DataDome and BotD respectively. We open source our rules for public use at this link.
Our key contributions are:
  • A large-scale analysis of browser fingerprint attributes for evasive bots that are able to evade anti-bot services.
  • A data-driven approach to discover inconsistencies in browser fingerprint attributes for detecting evasive bots.
@@
  • A novel use of honey site architecture to establish reliable ground-truth for evasive bots.
2 Background and Related Work
2.1 Evaluation of bot detection services
@@
Anti-bot services on the web generally generally employ machine learning to determine if an incoming request was sent by a human or a bot[21]. These services rely on several signals captured through different browser APIs, request headers, and behavior characteristics on a website[8]. Prior research has attempted to measure the accuracy of anti-bot services and understand their detection techniques. Azad et al. [18] analyzed 15 different anti-bot services, 14 of which used fingerprinting to detect bots. These services employ modern techniques such as WebGL and Canvas-based fingerprinting. They also detect inconsistencies in the collected bot fingerprints to determine if a request is from a bot or a human. We show in this paper (Section 7) that commercial anti-bot services can be more extensive in using inconsistencies to improve bot detection.
Azad et al. also tried to evaluate the performance of these services by deploying their own bots and measuring their evasiveness. They found that while most services were able to catch their ”Basic Bots”, more advanced bots, through either protection against fingerprinting scripts or employing fingerprints that are less common, were able to evade detection.
2.2 Analysis of bot traffic in the wild
@@
Xigao et. al [19] studied the prevalence of ”malicious” bots in the wild. They rely on the behavior of bots (indulging in credential stuffing, not honoring bots.txt etc) to characterize them as malicious. Such characterization is not applicable for bots indulging in impression fraud since these bots don’t display any specific behavior that can be leveraged for detection. One drawback of their approach is lack of a mechanism to entice evasive bots to visit their honey sites. While it is impossible for humans to accidentally land on their sites, the bots that visit the websites are also not likely to be evasive bots.
@@
Wu et. al[10] analyzed browser fingerprints from 36 billion requests on 14 commercial websites. Their analysis shows that adversarial bots (bots which change their fingerprints to avoid detection) have significantly different properties compared to benign bots. They found an overlap of only 1.6% of fingerprints between adversarial and benign bots. They also found that adversarial bots are involved in several different attack types, such as content scraping, fraudulent transactions, and credential stuffing. While Wu et. al conducted the largest study (at the time of writing) of bots in wild, their ground truth relies on decisions by F5 Inc.[22], a commercial anti-bot service. We show in this paper that evasive bots have high evasion rates against anti-bot services (Section 5), identifying the need to have a more robust mechanism to collect ground-truth. Next, we discuss some of the challenges affecting bot detection.
2.3 Challenges in bot detection
Detecting bots, especially those actively trying to evade detection, is a difficult task. Adversaries using bots employ various evasion techniques to bypass bot detection systems. We discuss some of the common evasion techniques employed by bots.
@@ 2.3 Challenges in bot detection
Polymorphism: Bots employ techniques such as morphing their User-Agent or other attributes[18] (i.e., fingerprints) to appear as benign website visitors. Prior research on bot detection has shown that morphing fingerprints can result in successful evasion. Iliou et. al [23] showed that while machine learning algorithms can detect simple bots with a precision and recall of 95% and 97% respectively, more advanced bots, i.e. bots that change their fingerprints, result in a drop in accuracy to only 55%.
@@
Behavioral Mimicry: Bots also simulate human-like behavior to evade behavioral analysis systems, including mimicking mouse movements, keystrokes, browsing patterns, and human text input[24]. Bot detection systems use these movements as “Human Interactive Proofs (HIPs)”[25, 26] to determine if a website visitor is a bot or a human. Jing et. al. [27] showed the development of a bot framework that employs behavioral mimicry. This framework enables bots to generate application-level events, such as keystrokes and mouse clicks, that closely resemble human actions to evade detection mechanisms. By analyzing inter-event timing and location information, the evasive bots successfully mimic human behaviors, posing a significant challenge for existing bot detection systems.
3 Threat Model
@@
In this paper, we specifically focus on bots committing impression fraud [3]. Web publishers who seek to artificially inflate the engagement on their websites indulge in this type of fraud. Inflating engagement allows these publishers to monetize and make profits on their websites through ads, even when they cannot guarantee visits to their website from legitimate users. Advertisers pay publishers for impressions of their ad (views, clicks etc) on the publisher’s website. However, only impressions recorded from legitimate users are useful to advertisers. Publishers who don’t receive traffic from legitimate users could employ bots to record these impressions to get paid by advertisers without delivering any useful impressions to them. We focus on bots indulging in impression fraud over other types of fraud (such as credential stuffing, account takeover etc), since these bots don’t have a need to perform specific actions to reach their goal, thereby making it more challenging to detect them.
@@
In our threat model, we consider publishers who incorporate anti-bot services on their websites to provide assurance of requests emerging from legitimate users, while employing evasive bots to evade detection.
4 Measurement infrastructure
@@
In this section, we describe our measurement infrastructure including the design of our novel honey site architecture. We design our measurement infrastructure to satisfy three requirements that enable us to reliably characterize evasive bots: first, we need reliable-ground truth that we only record requests from evasive bots of interest and no other entities (real users or other bots). Second, we need decisions from bot detection services on each request to identify isolate requests that evade detection. Third, we need to collect browser attributes that constitute browser fingerprints in these requests to analyze attributes that help with evasion.
4.1 Honey site architecture
@@
Figure 1: To collect requests from different bot services, we create multiple instances of the same honey site on the same domain. These instances differ in terms of random strings in their URL. We then drive traffic from different bot services to different instances of the honey site.
@@
The approach of using obscure domain names for honey sites  [19] cannot guarantee that the honey sites only receive requests from evasive bots. Bots that automatically send requests to such domains are typically indexing bots that visit new websites added to domain registries and other sources of DNS records. Examples of such bots includes search engine bots that do not have a need to conceal their identities. In fact, Google’s bots announce their identity through their User-Agent [28]. While evasive bots may also send requests such domains, the absence of a mechanism to isolate those requests makes it challenging to analyze them. evasive bots indulging in impression fraud do not have a need to perform any specific actions to reach their goal of recording views or impressions. Hence, such bots cannot be detected based on their actions/behavior.
@@
To overcome the challenge of only recording requests from evasive bots, we deploy multiple instances of the same honey site under the same domain. These versions only differ in terms of the presence of arbitrarily chosen strings in their URL but are identical in all other aspects. We do not record requests that do not contain one of these strings in the URL. This ensures that we do not record requests from real users or generic bots that discover our domain. We also share URLs with different arbitrary strings with different bot services. Thus, these URL strings also allow us to isolate requests received from different bot services. As a concrete example, if example.com is the domain of our honey site, example.com/XXXXX, example.com/YYYYY and example.com/ZZZZZ would constitute different instances of the honey site. We then purchase traffic from 3 different bot services to each send requests to one of these URLs. Real users and other generic bots who may stumble upon our site, will not know these strings and hence cannot send requests with these exact strings. Thus, by discarding requests that do not contain a valid URL string, we can ensure that we only record requests from bots we purchased. Figure 1 shows an overview of the honey site setup we use in our data collection.
4.2 Anti-bot services

### v2

@@ Abstract.
Browser fingerprinting is used for bot detection. In response, bots have started altering their fingerprints to evade detection. We conduct the first large-scale evaluation to study whether and how altering fingerprints helps bots evade detection. To systematically investigate such evasive bots, we deploy a honey site that includes two anti-bot services (DataDome and BotD) and solicit bot traffic from 20 different bot services that purport to sell “realistic and undetectable traffic.” Across half a million requests recorded on our honey site, we find an average evasion rate of 52.93% against DataDome and 44.56% evasion rate against BotD. Our analysis of fingerprint attributes of evasive bots shows that they indeed alter their fingerprints. Moreover, we find that the attributes of these altered fingerprints are often inconsistent with each other. We propose FP-Inconsistent, a data-driven approach to detect such inconsistencies across space (two attributes in a given browser fingerprint) and time (a single attribute at two different points in time). Our evaluation shows that our approach can reduce the evasion rate of evasive bots by 44.95%-48.11% while maintaining a true negative rate of 96.84% on traffic from real users.
1. Introduction
@@
The prevalence of bots on the web is on the rise (Erez Hasson, [n. d.]). Per recent reports, bots constitute around 49.6% of online traffic (imperva.com, [n. d.]), with 64.5% of those being bots that engage in malicious activity. Bad actors employ bots to launch a multitude of attacks (Springborn and Barford, 2013; Dave et al., 2013; Din et al., 2020; Nguyen Ba et al., 2021; Chiapponi et al., 2022). To counter such attacks, anti-bot services aim to detect and block bot traffic. Prior research has shown that anti-bot services use browser fingerprinting to detect bots without disrupting the user experience of legitimate users (Vastel et al., 2020; Asuman Senol and Bilogrevic, 2024). Browser fingerprints capture attributes of the web browser sending web requests and anti-bot services attempt to use differences in these attributes to distinguish bots from real users (Wu et al., 2023).
@@
Blackhat marketplaces (Babylon Traffic, [n. d.]; seoclerks, [n. d.]; Spark Traffic, [n. d.]), however, advertise realistic and undetectable bot traffic as a service. The traffic from such services constitute impression fraud and serve to artificially boost website engagement for monetization (Springborn and Barford, 2013; Farooqi et al., 2017; Javed et al., 2015). To evade detection, bots from these services are likely altering their fingerprint attributes that are used by anti-bot services for detection (Goßen et al., 2021; Jueckstock et al., 2021). We refer to such bots as evasive bots. It is important to characterize evasive bots and their fingerprints to improve the effectiveness of bot detection.
@@
Prior research has studied bot fingerprints by employing their own bots (Amin Azad et al., 2020; Vastel et al., 2020) or studying naturally discovered bots on their honey sites (Li et al., 2021). Thus, this line of work is not geared towards capturing the evasive fingerprints used by bots seeking to evade detection in the wild. Wu et al. performed a large-scale characterization of the differences between human and bot fingerprints in the wild (Wu et al., 2023). However, they did not specifically characterize evasive bots since they treat their bot detection system decisions as ground-truth to distinguish between the fingerprints of bots and real users.
@@
To fill this gap, we perform the first large-scale measurement of evasive bots that evade anti-bot services. To this end, we drive traffic from different bot services from blackhat marketplaces to different instances of our honey site. That way, the requests recorded at each honey site instance can be attributed to a bot service from whom we purchased traffic. These operators advertise their traffic as being realistic and natural, indicating that they likely employ evasive bots to ensure that they do not get detected. We integrate two commercial bot detection services (DataDome and BotD) on our honey site for bot detection. We also instrument our honey site to collect a range of fingerprint attributes.
@@
We collect 507,080 requests from 20 different bot services, with DataDome and BotD detecting 55.44% and 47.07% of these requests respectively. We analyze fingerprint attributes from different bot services to identify different sets of attributes that are effective at evading DataDome and BotD individually as well as attributes that are effective at evading both anti-bot services. Our analysis reveals spatial inconsistencies (among the different attributes of a given request) and temporal inconsistencies (across requests originating from the same device). These include obvious inconsistencies that cannot exist for real users, thereby making them useful signatures to detect bots.
@@
We use observations from our analysis to develop FP-Inconsistent, a data-driven approach to discover inconsistencies in fingerprint attributes for bot detection. FP-Inconsistent relies on the insight that real devices can only have a limited number of hardware and software configurations that are reflected in fingerprint attributes. Evasive bots, in their attempt to evade detection, emulate a large number of invalid or extraneous configurations. FP-Inconsistent leverages this mismatch between the expected and observed number of configurations to identify potential inconsistencies among evasive bot fingerprints. It does so by calculating the number of configurations for pairs of fingerprint attributes from evasive bots and identifying inconsistencies among attribute pairs that exhibit a higher-than-expected number of configurations.
Using this approach, we generate inconsistency rules that can be readily deployed by anti-bot services. Prior research focusing on the use of inconsistencies for bot detection (Vastel et al., 2018, 2020) has predominantly relied on one-off anecdotes to define inconsistencies that are not data-driven and hence do not scale. FP-Inconsistent systematizes the generation of inconsistency rules for bot detection.
Our evaluation shows the rules generated by FP-Inconsistent are able to achieve 48.11% and 44.95% reduction in traffic that evades DataDome and BotD respectively while maintaining a true negative rate of 96.84% on real user traffic. Our experiments also show that FP-Inconsistent does not incur false positives with most privacy-enhancing technologies. We open-source our honey site architecture and inconsistency rules for public use at this link.
Our work makes the following contributions:
@@
  • A novel honey site architecture to establish reliable ground-truth for evasive bots.
  • A large-scale measurement and analysis of fingerprint attributes for evasive bots that are able to evade anti-bot services.
  • A data-driven approach to discover inconsistencies in fingerprint attributes for detecting evasive bots.
2. Background and Related Work
2.1. Evaluation of bot detection services
@@
Anti-bot services on the web generally employ machine learning to determine if an incoming request was sent by a human or a bot(Cabri et al., 2018). These services rely on several signals captured through different browser fingerprinting APIs, request headers, and behavior characteristics on a website(Vastel et al., 2020). Prior research has attempted to measure the accuracy of anti-bot services and understand their detection techniques. Azad et al. (Amin Azad et al., 2020) analyzed 15 different anti-bot services, 14 of which used modern fingerprinting techniques such as WebGL and Canvas-based fingerprinting. While these services rely on inconsistencies in fingerprint attributes to detect bots, we show how they can be more extensive in using them to improve bot detection (Section 7).
Azad et al. also tried to evaluate the performance of these services by deploying their own bots and measuring their evasiveness. In contrast, we evaluate anti-bot services on requests from bots in the wild.
2.2. Analysis of bot traffic in the wild
@@
Xigao et al. (Li et al., 2021) studied the prevalence of ”malicious” bots in the wild. They use the behavior of bots (indulging in credential stuffing, not honoring bots.txt, etc) to characterize them as malicious. Such characterization is not applicable for bots indulging in impression fraud since these bots don’t exhibit any explicit behavior that can be leveraged for detection. Further, their approach draws traffic from bots in general and they do not include mechanisms to isolate evasive bots visiting their honey sites that seek to evade detection.
@@
Wu et al. (Wu et al., 2023) analyzed browser fingerprints from 36 billion requests on 14 commercial websites. Their analysis shows that adversarial bots (bots that change their fingerprints to avoid detection) have significantly different properties compared to benign bots. While they conducted the largest study (at the time of writing) of bots in the wild, their ground-truth relies on decisions by F5 Inc.(F5 Inc, [n. d.]), a commercial anti-bot service. Thus, without a more robust mechanism to collect ground-truth, their approach cannot analyze bots that can evade commercial anti-bot services such as F5.
Browser Polygraph (Kalantari et al., 2024) employs machine learning to detect bots that indulge in account takeover fraud (ATO). They predict if the fingerprint attributes in a request are consistent with the request’s reported User-Agent. In contrast, our work proposes a data-driven and semi-automatic technique to discover inconsistencies between any pair of fingerprint attributes (which includes but is not limited to the User-Agent) to combat impression fraud. Further, similar to the work of Wu et. al, their approach could be bolstered with more robust ground-truth since they rely on tags from FinOrg (a financial organization) to provide ground-truth for evaluation. In our work, our novel honey site architecture provides ground-truth to isolate traffic sent from different bot services.
2.3. Challenges in bot detection
We discuss some of the common techniques used by bots to evade detection.
@@ 2.3. Challenges in bot detection
Polymorphism: Certain bots morph their User-Agent or other attributes (i.e., fingerprints) to appear as benign website visitors for evasion (Kalantari et al., 2024; Amin Azad et al., 2020). Iliou et. al (Iliou et al., 2019) showed that while machine learning algorithms can detect simple bots with a precision and recall of 95% and 97% respectively, more advanced bots, i.e. bots that change their fingerprints, result in a drop in accuracy to only 55%.
@@
Behavioral Mimicry: Bots also simulate human-like behavior to evade behavioral analysis systems, including mimicking mouse movements, keystrokes, browsing patterns, and human text input(Askari et al., 2024). Bot detection systems use these movements as “Human Interactive Proofs (HIPs)”(Gianvecchio et al., 2008, 2009) to determine if a website visitor is a bot or a human. Jing et. al. (Jin et al., 2013) developed a bot framework for bots to generate keystrokes and mouse clicks that closely resemble human actions to evade detection.
3. Threat Model
@@
In this paper, we focus on bots committing impression fraud (Springborn and Barford, 2013). Web publishers who seek to artificially inflate the engagement on their websites indulge in this type of fraud. Inflating engagement allows these publishers to monetize and profit from their websites through ads, even when they cannot guarantee visits to their website from legitimate users. Advertisers pay publishers for impressions of their ad (views, clicks, etc) on the publisher’s website. However, only impressions recorded from legitimate users are useful to advertisers. Publishers who do not receive traffic from legitimate users could employ bots to record these impressions to get paid by advertisers without delivering any useful impressions to them. We focus on bots indulging in impression fraud over other types of fraud (such as credential stuffing, account takeover, etc), since these bots do not have a need to perform specific actions (Amin Azad et al., 2020; Li et al., 2021) to reach their goal, thereby making it more challenging to detect them.
@@
In our threat model, we consider publishers who incorporate anti-bot services on their websites to provide assurance of traffic from legitimate users, while employing evasive bots to evade detection.
4. Measurement infrastructure
@@
In this section, we describe our measurement infrastructure including the design of our novel honey site architecture. We design our measurement infrastructure to satisfy three requirements that enable us to reliably characterize evasive bots: first, we need reliable ground-truth that we only record requests from evasive bots of interest and no other entities (real users or other bots). Second, we need decisions from bot detection services on each request to isolate requests that evade detection. Third, we need to collect browser attributes that constitute browser fingerprints in these requests to analyze attributes that help with evasion.
4.1. Honey site architecture
@@
Figure 1. To collect requests from different bot services, we create multiple versions of the same honey site under the same domain. The only difference between these versions is the presence of different random strings in their URL. We then drive traffic from different bot services to different versions of the honey site.
@@
Using obscure domain names for honey sites  (Li et al., 2021) cannot guarantee that the honey sites only receive requests from evasive bots. Bots that automatically send requests to such domains are typically indexing bots that visit new websites added to domain registries and other sources of DNS records. Examples of such bots include search engine bots that do not have a need to conceal their identities. In fact, Google’s bots announce their identity through their User-Agent (Google, [n. d.]). While evasive bots may also send requests to such domains, the absence of a mechanism to isolate those requests makes it challenging to analyze them. Evasive bots indulging in impression fraud do not have a need to perform specific actions to record views or impressions. Hence, such bots cannot be detected based on their actions/behavior (Amin Azad et al., 2020; Li et al., 2021).
@@
To overcome the challenge of only recording requests from evasive bots, we deploy multiple versions of the same honey site under the same domain. These versions only differ in terms of the presence of arbitrarily chosen strings in their URL. We do not record requests that do not contain one of these strings in the URL to ensure that we do not record requests from real users or generic bots that discover our domain. We also share URLs having different arbitrary strings with different bot services. Thus, these URL strings enable the isolation of requests received from different bot services. As a concrete example, if example.com is the domain of our honey site, example.com/XXXXX, example.com/YYYYY, and example.com/ZZZZZ would constitute different versions of the honey site. We then purchase traffic from 3 different bot services to each send requests to one of these URLs. Real users and other generic bots who may stumble upon our site, will not know these strings and hence cannot include such strings in their requests. Thus, we can ensure that we only record requests from the bot services where we made our purchases using these URL strings. Figure 1 shows an overview of the honey site architecture.
4.2. Anti-bot services